OpenStudy (anonymous):

need help with substitutions please! 4x=3y-2 18=3x+y

OpenStudy (anonymous):

sorry - this shoudnt be here. from second equation y=18-3x substitute this into first equation: 4x = 3(18-3x) -2 solve this for x and plug this value of x into the second equation to get value of y

OpenStudy (anonymous):

So is that now 4x=54-9x-2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

13x = 52 x=4

OpenStudy (anonymous):

I got it now, I don't know what I was thinkin!

OpenStudy (anonymous):

no probs - plz click 'good answer' to close question - if u need any more help plz post

OpenStudy (anonymous):

thanx